# Gas price shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v1

@@ 4 Conclusions
This paper provides new evidence on how natural gas price shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. Using quarterly survey data from the Bank of Italy’s SIGE spanning over two decades, we identify structural natural gas price shocks through a Bayesian VAR framework with sign and zero restrictions. Our analysis reveals three main findings. First, natural gas price shocks are a significant driver of firms’ inflation expectations throughout the sample period, with particularly strong effects during the post-COVID energy crisis of 2021-2023. Historical decomposition shows that these shocks predominantly exerted downward pressure on expectations during normal times but contributed substantially to the surge in inflation expectations following Russia’s invasion of Ukraine. Second, firms respond to natural gas price innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock. Both aggregate inflation and firms’ own prices exhibit similar dynamics, with effects dissipating within 5-6 quarters, consistent with the transitory nature of energy price shocks documented in the literature. Inflation uncertainty also increases in response to the shock, suggesting that energy price disruptions affect not only the level of prices but also firms’ confidence in forecasting future inflation. Third, and most importantly, we document substantial non-linearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weak demand conditions prevent cost pass-through.

### v3

@@ 4 Conclusions
This paper provides new evidence on how natural gas supply shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. We identify a gas supply shock through an external IV-VAR framework and we analyze its implications for firms’ pricing using quarterly survey data from the Bank of Italy’s SIGE, spanning over two decades. Our analysis reveals three main findings. First, natural gas supply shocks emerge as a significant driver of electricity prices in Italy. An 1% increase in the spot natural gas price measured by the Dutch TTF future determines a 0.5% increase in the Italian electricity price, measured by the PUN on impact. Second, firms respond to natural gas supply innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock, with effects dissipating within 5-6 quarters, consistent with the transitory nature of energy price shocks documented in the literature. The increase of inflation uncertainty in response to the shock suggests that energy price disruptions affect not only the level of prices but also firms’ ability in forecasting future inflation. Third, and most importantly, we document substantial non-nonlinearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weak demand conditions prevent cost pass-through.